The market research Department of Paradox Enterprises has determined that the demand for fingolds is ๐‘„ = 1,000 โˆ’ 5๐‘ƒ + 0.05๐‘€ โˆ’ 50๐‘ƒ๐‘ง where P is the price of glibdips, M is income, and ๐‘ƒ๐‘ง is the price of ballzacks. Suppose that P = $5, M = $20,000, and ๐‘ƒ๐‘ง = $15. a. Calculate the price elasticity of demand for fingolds (3mks) b. Is the firm maximizing its total revenue at P = $5. If not, what price should it charge? (3mks) c. At P = $5, compute the income elasticity of demand for fingolds(3mks) d. At P = $5, cross-price elasticity of demand for fingolds.ย 


1
Expert's answer
2022-01-31T09:57:33-0500

"Q=1000-5P+0.05M-50P_z"

"P" =$5

"M" =$20000

"P_z=" $15

"1000-5(5)+0.05(20000)-50(15)"

"1000-25+1000-750=1225"

a)"E_p=\frac{\Delta Q}{\Delta P} \times \frac{P}{Q}"

"-5\times\frac{5}{1225}=-0.02" , inelastic.

b) "R=Q\times P=1225\times 5=6125"

Yes, the firm is maximizing its total revenue at this price.

c)"E_i=\frac{dQ}{dI}\times\frac{I}{X}=0.05\times\frac{20000}{5}=200"

d)"E_cp=\frac{dQ}{dP}\times \frac{P}{Q}=-50\times\frac{5}{1225}=-0.2" , inelastic.